Atlanta Cement, Inc. buys on terms of 2/15, net 30. It does not take discounts, and it typically pays 115 days after the invoice date. Net purchases amount to $720,000per year. What is the nominal annual percentage cost of its non-free trade credit, based on a 365-day year?
Answer #1
Many industries or trades will have typical terms, such as “2/15 net 30.” This means the entire balance is due within 30 days of the invoice, but the buyercan take a 2% cash discount if the bill is paid within 15 days. Buyers should always take cash discounts, even if they have to borrow money to do so(unless the loan interest rate is very high).

Nominal % Cost= Disc % / (100-disc %) X (365/ act. Days – Disc Days)

(2%)/ (100-2)(365/115-15) = (0.02)/(0.98)(3.65)
= 0.02 / 3.577
= 0.00559
= 0.5.59%
answered by: brenda smith